# Intertwined superconductivity and orbital selectivity in a three-orbital Hubbard model for the iron pnictides

### v1

@@ Abstract
We finally notice the role of the underlying band structure. Indeed, a simpler tight-binding Hamiltonian $\mathcal{H}_{0}$ with the same diagonal nearest-neighbor hopping for every orbital, displays superconductivity for $n=3+1/3$, but with the $d_{x^{2}-y^{2}}$ symmetry sup , that one obtains for the one-band Hubbard model Tocchio et al. (2016). This finding underlies that, even in a model with only local electronic interactions, the symmetry of the order parameter depends on the shape of the Fermi surface.
@@
Conclusions. By using variational wave functions including pairing parameters and Jastrow factors, we have studied a simplified model for IBS superconductors. Spanning the density range between $n=3$ (Mott insulator) and $n=4$ (Hund’s metal), we find a nematic state with orbital-selective character and superconducting pairing. In particular, the $d_{yz}$ (or $d_{xz}$) orbital remains half-filled in the whole density regime, thus leading to a spontaneous orbital-selective Mott transition and a broken rotational symmetry. Most importantly, superconductivity is observed in the nematic regime, changing its character when the electron doping is varied. For $3<n\lesssim 3.5$, the $d_{xy}$ orbital sustains long-range electron pairing. Instead, for $43.5\lesssim n<$ 4, finite pairing correlations are observed in the $d_{xz}$ (or $d_{yz}$) orbital, even though the BCS parameters are small in the superconducting orbital; in this unconventional case, electron pairing can be established only through the hybridization with the other orbitals. All these features are triggered by the presence of Hund’s coupling, since neither orbital selectivity nor superconductivity are found for $J=0$. The tentative phase diagram for $3\leq n\leq 4$ is shown in Fig. 4.
@@
We acknowledge useful discussions with L. Fanfarillo. Computational resources were provided by HPC@POLITO (http://www.hpc.polito.it). We acknowledge also the CINECA award under the ISCRA initiative, for the availability of high performance computing resources and support. M.C. acknowledges financial support of MUR via PRIN 2020 (Prot. 2020JLZ52N 002) programs, and by the European Union - NextGenerationEU through PRIN 2022 (Prot. 20228YCYY7), National Recovery and Resilience Plan (NRRP) MUR Project No. PE0000023-NQSTI, and No. CN00000013-ICSC. The views and opinions expressed are solely those of the authors and do not necessarily reflect those of the European Union, nor can the European Union be held responsible for them.
@@ I.1 Three-orbital Hubbard model
I.2 Fourier transform of the BCS parameters at density $n=3+1/3$
@@
I.3 Results for the simplified multiorbital model

### v2

@@ Abstract
We finally notice that, even for purely local interactions, the symmetry of the order parameter depends on the shape of the Fermi surface. Indeed a simpler tight-binding Hamiltonian $\mathcal{H}_{0}$ with the same diagonal nearest-neighbor hopping for every orbital, displays superconductivity for $n=3+1/3$, but with the $d_{x^{2}-y^{2}}$ symmetry sup , that one obtains for the one-band Hubbard model Tocchio et al. (2016).
@@
Conclusions. By using variational wave functions including pairing parameters and Jastrow factors, we have studied a simplified model for IBS superconductors. Spanning the density range between $n=4$ (parent compound) and $n=3$ (Mott insulator), we find a nematic state with orbital-selective character and superconducting pairing. In particular, the $d_{yz}$ (or $d_{xz}$) orbital remains half-filled in the whole density regime, thus leading to a spontaneous orbital-selective Mott transition and a broken rotational symmetry. Most importantly, superconductivity is observed in the nematic regime, changing its character when the electron doping is varied. For $3<n\lesssim 3.5$, the $d_{xy}$ orbital sustains long-range electron pairing. Instead, for $43.5\lesssim n<$ 4, pairing correlations are dominant in the $d_{xz}$ (or $d_{yz}$) orbital. All these features are triggered by the presence of Hund’s coupling, since neither orbital selectivity nor superconductivity are found for $J=0$. The tentative phase diagram for $3\leq n\leq 4$ is shown in Fig. 4.
@@
Acknowledgements. We acknowledge useful discussions with L. Fanfarillo. Computational resources were provided by HPC@POLITO (http://www.hpc.polito.it). We acknowledge also the CINECA award under the ISCRA initiative, for the availability of high performance computing resources and support. M.C. acknowledges financial support of MUR via PRIN 2020 (Prot. 2020JLZ52N 002) programs, and by the European Union - NextGenerationEU through PRIN 2022 (Prot. 20228YCYY7), PNRR MUR Projects No. PE0000023-NQSTI, and No. CN00000013-ICSC. The views and opinions expressed are solely those of the authors and do not necessarily reflect those of the European Union, nor can the European Union be held responsible for them.
@@ I Supplemental Material
In this Supplemental Material, we show the tight-binding model ${\cal H}_{0}$ that is used for the three-orbital Hubbard Hamiltonian considered in this work. Then, we present a comparison between the intra-orbital pairing correlations $D_{\alpha}(r)$ at density $n=3+1/3$ and $n=3+2/3$, within the orbital selective region and at $U=0$. Moreover, we show the optimal BCS parameters as a function of $U$ at density $n=3+1/3$ and $n=3+2/3$, as well as the symmetry of the BCS parameters in $k$-space at density $n=3+1/3$. Finally, we present results for the BCS parameters and for the pairing correlations when the tight-binding Hamiltonian is studied in a simplified version with only nearest-neighbor intra-orbital hoppings (all equal for the three orbitals).
@@ I.1 Three-orbital Hubbard model
Figure 6: Upper panels: $D_{\alpha}(r)$ in all orbitals, for a value of $U$ within the orbital-selective regime. Data are shown at total electronic densities $n=3+1/3$ (left column) and $n=3+2/3$ (right column). Lower panels: The same as above, but for $U=0$. Pairing correlations are obtained on a $L=18\times 18$ cluster at $J/U=0.2$.
Figure 7: Optimal intra-orbital BCS parameters [see Eq. (2) of the main text] as a function of $U$. Data are shown at total electronic densities $n=3+1/3$ (left column) and $n=3+2/3$ (right column). BCS parameters are computed with the variational approach on a $L=12\times 12$ lattice at $J/U=0.2$.
@@
I.2 Pairing correlations at $U=0$
In Fig. 6, we report the intra-orbital pairing correlations $D_{\alpha}(r)$ at densities $n=3+1/3$ and $n=3+2/3$, both within the orbital-selective phase and at $U=0$. In the former case, we observe that pairing correlations at large distance are clearly enhanced going from $U=0$ to $U=2$eV. On the contrary, in $d_{xz}$ and $d_{yz}$ orbitals, they are suppressed when electron-electron correlation is switched on. These features clearly indicate that only the $d_{xy}$ orbital becomes superconductive. At density $n=3+2/3$, the situation is completely different. Pairing correlations at large distance are strongly enhanced in the $d_{xz}$ orbital when moving from $U=0$ to $U=2.3$eV, while they remain similar or are even suppressed in the other two orbitals when electron-electron correlation is present. In this case, superconductivity develops in the $d_{xz}$ orbital.
I.3 Optimal BCS parameters
Within the VMC approach, the presence of superconductivity is usually associated with the stabilization of BCS parameters in the optimized wave function, see Eq. (2) in the main text, that defines the uncorrelated state |$\ket{\Phi_{0}}$. Still, pairing in the uncorrelated part of the wave function does not imply superconducting order, since the Jastrow factor may modify or even suppress the actual pairing correlations (e.g., in the Mott regime Capello et al. (2005); De Franco et al. (2018)).
In Fig. 7, we show the optimal (intra-orbital) BCS parameters $\Delta_{\alpha,\delta}$ of Eq. (2) of the main text, for $n=3+1/3$ and $n=3+2/3$. In both cases, at sufficiently large values of the Hubbard-$U$ interaction, the orbital selective phase is accompanied by the development of sizable BCS parameters in both $d_{xy}$ and $d_{yz}$ orbitals, while they are relatively small in the $d_{xz}$ one. However, the presence of inter-orbital hoppings in the Hamiltonian, in addition to the Jastrow factors, induces a remarkable difference in the pairing correlations: For $n=3+1/3$, the stabilization of finite BCS parameters in the $d_{xy}$ orbital leads to a true superconducting (long-range) correlation $D_{xy}(r)$, see Fig. 6. Instead, the other two orbitals, whose densities are close to half-filling, show a very small signal (in this case, the Jastrow factor gives a substantial suppression of pairing correlations). The situation is different for $n=3+2/3$, where $n_{xy}\approx 1$ and the pairing correlations $D_{xy}(r)$ are much reduced (again by the Jastrow factor) with respect to the previous case, see Fig. 6 ($U=2.3$eV). However, surprisingly, the $d_{xz}$ orbital (which is not half-filled) shows sizable pairing correlations, despite having relatively small BCS parameters. This fact is only possible because of the inter-orbital hybridization in the BCS Hamiltonian.
I.4 Fourier transform of the BCS parameters at density $n=3+1/3$
@@
I.5 Results for the simplified multiorbital model